A stone is dropped into water from a bridge 44.1 m above the water. Another stone is thrown vertically downward 1 sec later. Both strike speed water simultaneously. What was the initial speed of the second stone?


                (a)           12.25/s


                (b)           14.75m/s


                (c)           16.23m/s


                (d)           17.15m/s

Grade:10

2 Answers

Keerthi Raj
10 years ago

placing origin of the coordinate system at the point of dropping.

-44.1=-4.9t2 ==> t=3s

-44.1=2u-19.6 ==> u=-12.25m/s

Yash Deshpande
5 years ago
x=44.1m
a=9.8m/s2
u=0
Wkt,
x=ut+1/2gt2
44.1=1/2*9.8*t2
t2=441/49
t=3 sec. So the second stone will come down in 3-1=2sec
Therefore, applying the same eqn again we get,
44.1=2u+1/2*9.8*4
44.1=2u+19.6
u=12.25m/s